# Module-conditioned distribution of quantum circuits

### v2

@@ 4.1.1 $CZ$ fraction circuits
8:        With probability $1-p$, add a unary gate acting on $q_{i}$ to $U$.
@@
Figure 8: An example $7$-qubit $3$-layer $CZ$ fraction circuit with $p=4/7$.
Figure 9: BIP improvements observed in $CZ$ fraction circuits. The reduction in ebit cost generally increases with higher $CZ$ fractions, since they correspond to more difficult DQC instances.
Given parameters $(n,d,p)$, an $n$-qubit $CZ$ fraction circuit is generated as described in Algorithm 3 [17]. Fig. 8 shows an example $CZ$ fraction circuit. In all plots, ‘HP’ stands for the hypergraph partitioning solution. Ebit costs for distributing random $CZ$ fraction circuits are shown in Fig. 9.
@@ 4.1.2 Quantum volume circuits
Figure 10: An example $6$-qubit $2$-layer quantum volume circuit before transpilation.
Figure 11: BIP improvements observed in quantum volume circuits. The reduction in ebit cost is relatively marginal for these circuits. This is because as mentioned in [1], rewriting a given quantum volume circuit such that all non-unary gates are CP gates makes the fraction of binary gates quite low. Similar to the results for circuits with low $CZ\_\text{percent}$ shown in Fig. 9, little improvement is expected.
Given parameters $(n,d)$, an $n$-qubit quantum volume circuit is generated as described in Algorithm 4 [11]. Fig. 10 (right) shows an example quantum volume circuit. Ebit costs for distributing random quantum volume circuits are shown in Fig. 11.
@@ 4.2.1 QFT circuits
Ebit costs for distributing QFT circuits are shown in Fig. 12. We remark that replacing $\pi$ (found by a hypergraph partitioner) with $\pi^{*}$ defined in (10) improved the ebit cost in every single run of the BIP solver. In this regard, we conjecture that for the distribution of QFT circuits, $\pi^{*}$ is not only the optimal module allocation function for MS-HC but also for MS-GC.
Figure 12: BIP and/or $\pi^{*}$ improvements observed in QFT circuits.
@@ 4.2.2 DraperQFTAdder
The DraperQFTAdder is a quantum circuit that leverages QFT to perform addition [13]. Before distribution, we perform a minor optimization on this circuit. Consider the identity
@@
Figure 13: (a) Example DraperQFTAdder circuit that performs in-place addition (modulo $2^{2}$) on two $2$-qubit registers. (b) Circuit representation of (12). (c) Equivalent circuit obtained by applying (12).
Figure 14: BIP improvements observed in DraperQFTAdder circuits.
where the subscripts denote qubit indices (see APPENDIX C). The right-hand side appears many times in DraperQFTAdder circuits (Fig. 13). Since SWAP gates require binary gates to implement, we replace the right-hand side of (12) with the left-hand side. Ebit costs for distributing DraperQFTAdder circuits are shown in Fig. 14.
@@ 4.2.3 RGQFTMultiplier
The RGQFTMultiplier is a quantum circuit that leverages QFT to perform multiplication [27]. They store the product of two $n^{\prime}$-bit inputs out-of-place. By default, the output register has $2n^{\prime}$ qubits, resulting in a total of $4n^{\prime}$ qubits (Fig. 15).
Figure 16: BIP improvements observed in RGQFTMultiplier circuits.
Ebit costs for distributing RGQFTMultiplier circuits are shown in Fig. 16. Surprisingly, the hypergraph partitioner alone returns extremely inefficient distributions in many cases; post-processing with the BIP solver reduces the ebit costs significantly.
@@ 4.3.1 AND circuits
Figure 17: (a) Example logical AND operation on 3 qubits. (b) Example $4$-qubit Boolean function that computes the inner product of two $2$-qubit vectors over $F_{2}$.
Figure 18: BIP improvements observed in AND circuits.
The AND circuit implements the logical AND operation on a number of qubits, i.e., it is a multi-controlled $X$ gate (Fig. 17a). Ebit costs for distributing AND circuits are shown in Fig. 18.
@@ 4.3.2 InnerProduct circuits
Figure 19: BIP improvements observed in InnerProduct circuits. Empty bars indicate zero ebit cost.
@@
$\mathcal{IP}_{2n}(x_{1},\cdots,x_{n},y_{1},\cdots,y_{n})=(-1)^{x\cdot y}.$
Ebit costs for distributing InnerProduct circuits are shown in Fig. 19.
5 Conclusion and outlook
In this work, we proposed BIP formulations for the distribution of quantum circuits with fixed module allocation function. We take the module allocation returned by the hypergraph partitioning formulation of DQC, which had previously achieved the best results for $k\geq 3$ modules. Combining our post-processing step with existing methods for module allocation yields new best results. We also show that the problem can be further optimized for $k=3$ by leveraging the unique properties of this specific case.
As is typical for DQC, our method is applicable to any quantum circuits constructed exclusively with unary gates and $CP$ gates. Of these, QFT circuits have been found to reveal special properties in experiments. That is, taking the conventional ordering of qubits as the module allocation is optimal for MS-HC. Based on numerical results, we further conjecture that this module allocation is also optimal for MS-GC.
We believe that for many practical circuits, observing local properties may enable further reduction in the number of BIP variables and/or constraints. Identifying these properties is a promising future research direction to enhance the scalability of our approach. We also look forward to developing integer programming models for other circuit distribution protocols beyond the cat-entanglement-based approach and/or addressing connectivity between modules [15].
Appendix A Cat-entanglements commute with $CP$ gates
Figure 20: A linked copy $e$ of $q$ is created at the module containing $q^{\prime}$. The order of cat-entanglement and $CP$ does not matter.
Appendix B Cat-entanglements do not commute with unary gates
@@
Appendix C Derivation of (12)
In matrix representations,

### v3

@@ 4.1.1 $CZ$ fraction circuits
8:   With probability 1p1-p, add a unary gate acting on $q_{i}$ to $U$.
@@
Figure 9: An example 77-qubit 33-layer CZCZ fraction circuit with $p=4/7$.
Given parameters $(n,d,p)$, an $n$-qubit CZCZ fraction circuit is generated as described in Algorithm 3 [g2021efficient]. Fig. 9 shows an example $CZ$ fraction circuit. It is observed from Table 3 that as the $CZ$ fraction increases from 10% to 90%, the maximum ebit saving achieved by the BIP post-processing grows. Ebit costs for distributing random $CZ$ fraction circuits are shown in Fig. 16 in Appendix D.1.
@@ 4.1.2 Quantum volume circuits
Figure 10: An example 66-qubit 22-layer quantum volume circuit before transpilation.
Given parameters $(n,d)$, an $n$-qubit quantum volume circuit is generated as described in Algorithm 4 [cross2019validating]. Fig. 10 (right) shows an example quantum volume circuit. Ebit costs for distributing random quantum volume circuits are shown in Fig. 17 in Appendix D.2.
@@ 4.2.1 QFT circuits
Ebit costs for distributing QFT circuits are shown in Fig. 18 in Appendix D.3. We remark that replacing $\pi$ (found by a hypergraph partitioner) with $\pi^{*}$ defined in Eq. (12) improved the ebit cost in every single run of the BIP solver. In this regard, we conjecture that for the distribution of QFT circuits, $\pi^{*}$ is not only the optimal module allocation function for MS-HC but also for MS-GC.
@@ 4.2.2 DraperQFTAdder
Figure 11: (a) Example DraperQFTAdder circuit that performs in-place addition (modulo $2^{2}$) on two 22-qubit registers. (b) Circuit representation of Eq. (14). (c) Equivalent circuit obtained by applying Eq. (14).
The DraperQFTAdder is a quantum circuit that leverages QFT to perform addition [draper2000addition]. Before distribution, we perform a minor optimization on this circuit. Consider the identity
@@
where the subscripts denote qubit indices (see Appendix D.4). The right-hand side appears many times in DraperQFTAdder circuits (Fig. 11). Since SWAP gates require binary gates to implement, we replace the right-hand side of Eq. (14) with the left-hand side. Ebit costs for distributing DraperQFTAdder circuits are shown in Fig. 19 in Appendix D.4.
@@ 4.2.3 RGQFTMultiplier
The RGQFTMultiplier is a quantum circuit that leverages QFT to perform multiplication [ruiz2017quantum]. It stores the product of two $n^{\prime}$-bit inputs out-of-place. By default, the output register has $2n^{\prime}$ qubits, resulting in a total of $4n^{\prime}$ qubits (Fig. 12).
Ebit costs for distributing RGQFTMultiplier circuits are shown in Fig. 20 in Appendix D.5. Surprisingly, the hypergraph partitioner alone returns extremely inefficient distributions in many cases, and BIP post-processing reduces the ebit costs significantly.
@@ 4.3.1 AND circuits
Figure 13: (a) Example logical AND operation on 3 qubits. (b) Example 44-qubit Boolean function that computes the inner product of two 22-qubit vectors over $F_{2}$.
The AND circuit implements the logical AND operation on a number of qubits, i.e., it is a multi-controlled-$X$ gate (Fig. 13a). Ebit costs for distributing AND circuits are shown in Fig. 21 in Appendix D.6.
@@ 4.3.2 InnerProduct circuits
InnerProduct is a 2nn-qubit Boolean function that computes the inner product of two $n$-qubit vectors over $F_{2}$ (Fig. 13b):
$\mathcal{IP}_{2n}(x_{1},\cdots,x_{n},y_{1},\cdots,y_{n})=(-1)^{x\cdot y}.$
Ebit costs for distributing InnerProduct circuits are shown in Fig. 22 in Appendix D.7.
5 Conclusion and outlook
We proposed exact BIP formulations for the DQC problem. Given a fixed module allocation function, we optimize the remaining migration-selection layer to minimize ebit cost. This decoupling turns the fixed-allocation subproblem into a step that can be used as post-processing on top of any allocation method, including hypergraph partitioning. We further derived a tighter specialization for the three-module case, leveraging the fact that it admits a reduced migration search space.
Our experiments indicate that BIP post-processing is a reliable choice for distributions obtained from heuristic allocators. Concretely, we observed reductions of up to 20% for random circuits, and in several structured arithmetic benchmarks the reduction exceeded an order of magnitude. We also found that the magnitude of improvement depends on circuit structure. For instance, circuits with higher fractions of non-local gates tend to admit larger savings.
The BIP is performed offline during compilation and therefore introduces additional classical runtime, but this overhead can be amortized when the same distributed circuit is executed arbitrarily many times. In this regime, even moderate reductions in ebit cost translate directly into recurring savings in entanglement consumption and classical communication overhead at execution time.
Several directions can further improve scalability and broaden applicability. First, many practical circuits exhibit local regularities (e.g., repeated gate motifs or restricted interaction patterns). Exploiting such structure to eliminate redundant variables and constraints could substantially speed up the optimization. Second, while we highlighted quantum Fourier transform as a structured case where allocation admits principled choices, there may be other structured families where the optimal allocation can be nicely characterized. Combining such allocation rules with BIP post-processing is a promising avenue. Finally, it would be valuable to extend integer-programming models beyond the cat-entanglement-based protocol, and to incorporate additional system constraints such as restricted inter-module connectivity or heterogeneous link costs [ferrari2023modular].
Appendix A Implementing non-local controlled unitary gate using a shared Bell pair
This section details the underlying mechanics of Fig. 2. Let $a,e,e^{\prime},b$ be single-qubit registers. Let $U$ be an arbitrary single-qubit unitary acting on $b$. We assume the input on $ab$ is an arbitrary pure state
@@ Appendix A Implementing non-local controlled unitary gate using a shared Bell pair
and $ee^{\prime}$ share the Bell pair
[…]
We show that the following procedure implements controlled-$U$ with control $a$ and target $b$:
  1. 1.
    Apply $\mathrm{CNOT}$ with control $a$ and target $e$.
  2. 2.
    Measure $e$ in the computational basis, obtaining outcome $m\in\{0,1\}$. If $m=1$, apply $X$ on $e^{\prime}$.
  3. 3.
    Apply controlled-$U$ with control $e^{\prime}$ and target $b$.
  4. 4.
    Apply $H$ on $e^{\prime}$, measure $e^{\prime}$ in the computational basis, obtaining outcome $n\in\{0,1\}$. If $n=1$, apply $Z$ on $a$.
The register order is permuted as needed to maintain readability.
Step 0 (initial state). The joint initial state on $a,e,e^{\prime},b$ is
[…]
Step 1 (apply $\mathrm{CNOT}_{a\to e}$). Applying $\mathrm{CNOT}$ with control $a$ and target $e$ flips $e$ iff $a=1$. A direct expansion gives
[…]
Grouping by the computational basis of $e$ yields
[…]
Step 2 (measure $e$ and conditionally correct $e^{\prime}$). Measure $e$ in the computational basis, obtaining outcome $m\in\{0,1\}$. If $m=0$, the (unnormalized) post-measurement state on $a,e^{\prime},b$ is
[…]
If $m=1$, the (unnormalized) post-measurement state on $a,e^{\prime},b$ is
[…]
Therefore, after the classical feed-forward correction, independently of $m$ the (unnormalized) state can be written as
[…]
In particular, $e^{\prime}$ is perfectly correlated with $a$ in the computational basis.
Step 3 (apply controlled-$U$ with control $e^{\prime}$ and target $b$). Applying $CU_{e^{\prime}\to b}$ to Eq. (A) yields
[…]
Now measure $e^{\prime}$ in the computational basis, obtaining outcome $n\in\{0,1\}$. Conditioned on $n$, the (unnormalized) post-measurement state on $ab$ is:
[…]
Step 5 (apply $Z$ on $a$ if $n=1$). If $n=1$, apply $Z$ on $a$. Since $|0Z\ket$$\ket{0}$ and $|1Z\ket$$\ket{1}$, this removes the relative minus sign in Eq. (18), giving
[…]
independently of the measurement outcomes $m,n$.
Identify $CU_{a\to b}$. Writing the original input Eq. (15) as
[…]
we see that Eq. (19) is exactly the result of applying controlled-$U$ with control $a$ and target $b$, i.e.,
[…]
Hence the protocol implements the desired non-local controlled-$U$ gate.
Appendix B Cat-entanglements commute with $CP$ gates
Figure 14: A linked copy $e$ of $q$ is created at the module containing $q^{\prime}$. The order of cat-entanglement and CPCP does not matter.
Appendix C Cat-entanglements do not commute with unary gates
@@
Appendix D Additional experimental results
This section provides the full ebit-cost plots for all benchmark families in Section 4. Additionally, the execution time for each typical BIP instance is shown as a line plot overlaid on the corresponding bars.
D.1 $CZ$ fraction circuits
Figure 16: BIP improvements observed in CZCZ fraction circuits. Here, $n$ is the number of qubits, $d$ is the circuit depth, and $k$ is the number of modules. The reduction in ebit cost generally increases with higher CZCZ fractions, since they correspond to more difficult DQC instances.
D.2 Quantum Volume circuits
Figure 17: BIP improvements observed in quantum volume circuits. Here, $n$ is the number of qubits and $k$ is the number of modules. The reduction in ebit cost is relatively marginal for these circuits. This is because as mentioned in [andres2024distributing], rewriting a given quantum volume circuit such that all non-unary gates are CPCP gates makes the fraction of binary gates quite low. Similar to the results for circuits with low $CZ\_\text{percent}$ shown in Fig. 16, little improvement is expected.
D.3 QFT circuits
Figure 18: BIP and/or $\pi^{*}$ improvements observed in QFT circuits. Here, $k$ is the number of modules and $m$ is the number of qubits per module. Applying BIP to $\pi^{*}$ consistently yields better results than applying it to an HP-based allocation. This demonstrates that the current HP is suboptimal not only in how it selects migrations but also in the module allocations it produces.
D.4 DraperQFTAdder circuits
We first derive Eq. (14). In matrix representations,
@@ D.4 DraperQFTAdder circuits
Figure 19: BIP improvements observed in DraperQFTAdder circuits. Here, $n$ is the number of qubits and $k$ is the number of modules.
D.5 RGQFTMultiplier circuits
Figure 20: BIP improvements observed in RGQFTMultiplier circuits. Here, $n$ is the number of qubits and $k$ is the number of modules. In many configurations, our BIP post-processing achieves significant ebit savings compared to HP.
D.6 AND circuits
Figure 21: BIP improvements observed in AND circuits. Here, $n$ is the number of qubits and $k$ is the number of modules. In many configurations, our BIP post-processing achieves significant ebit savings compared to HP.
D.7 InnerProduct circuits
Figure 22: BIP improvements observed in InnerProduct circuits. Here, $n$ is the number of qubits and $k$ is the number of modules. Empty bars indicate zero ebit cost. In many configurations, our BIP post-processing achieves significant ebit savings compared to HP.
Author contributions
H.C. provided the main ideas, conducted the experiments, and wrote the manuscript. J.L. analyzed the results and revised the manuscript.
Data availability
No datasets were generated or analyzed during the current study.
Declarations
Competing interests
The authors declare no competing interests.